# Having Beer after Prayer? Measuring Cultural Bias in Large Language Models

### 19 Feb 2024 v3

@@ Abstract
As the reach of large language models (LMs) expands globally, their ability to cater to diverse cultural contexts becomes crucial. Despite advancements in multilingual capabilities, models are not designed with appropriate cultural nuances. In this paper, we show that multilingual and Arabic monolingual LMs exhibit bias towards entities associated with Western culture. We introduce CAMeL, a novel resource of 628 naturally-occurring prompts and 20,368 entities spanning eight types that contrast Arab and Western cultures. CAMeL provides a foundation for measuring cultural biases in LMs through both extrinsic and intrinsic evaluations. Using CAMeL, we examine the cross-cultural performance in Arabic of 12 different LMs on tasks such as story generation, NER, and sentiment analysis, where we find concerning cases of stereotyping and cultural unfairness. We further test their text-infilling performance, revealing the incapability of appropriate adaptation to Arab cultural contexts. Finally, we analyze 6 Arabic pre-training corpora and find that commonly used sources such as Wikipedia may not be best suited to build culturally aware LMs, if used as they are without adjustment. We will make CAMeL publicly available at: https://github.com/tareknaous/camel
@@
We show that CAMeL entities and prompts enable cross-cultural testing of LMs in versatile experimental setups, including story generation, NER, sentiment analysis, and text infilling (§4). We benchmark 12 LMs pre-trained with Arabic data (§4.1). Our results reveal concerning cases of cultural stereotypes in LM-generated stories, such as the association of Arab names with poverty/traditionalism (§4.2), and cultural unfairness, such as better NER tagging performance of Western entities and higher association of Arab entities with negative sentiment (§4.3). We further show that LMs exhibit high levels of preference towards Western-associated entities even when prompted by contexts uniquely suited for Arab culture-associated entities (§4.4).
@@ 4.1 Language Models
We consider LMs that have been intentionally trained for Arabic. For monolingual LMs, we use AraBERT Antoun et al. (2020) and ARBERT Abdul-Mageed et al. (2021). We also include MARBERT Abdul-Mageed et al. (2021) and AraBERT-T Antoun et al. (2020), which are trained on Arabic tweets, as well as AraGPT2 Antoun et al. (2021). For multilingual LMs, besides mBERT, XLM-R Conneau et al. (2020), BLOOM Scao et al. (2022), GPT-3.5 and GPT-4, we use Arabic-English bilingual JAIS Sengupta et al. (2023), GigaBERT and GigaBERT-CS Lan et al. (2020), which was further trained on synthetic Arabic-English code-switched data. We use the base () and large () versions whenever available.
@@ 4.3 Fairness in NER and Sentiment Analysis
We create models capable of performing Arabic NER and sentiment prediction by fine-tuning BERT-type LMs on datasets commonly used in Arabic NLU benchmarks Elmadany et al. (2023); Abdul-Mageed et al. (2021). We use the ANERCorp Benajiba et al. (2007) dataset for NER (name and location tags were used only) and HARD dataset Elnagar et al. (2018) for sentiment analysis. For GPT-type LMs, we perform in-context learning with 5-shot examples (see prompts in Appendix F.2).
@@ 4.4 Culturally-Appropriate Text Infilling
where is the LM’s probability of an entity filling the masked token. We evaluate LMs with BERT-type architecture using the full prompts with a [MASK] token for text-infilling and LMs with GPT-type architecture using only the portion of the prompt appearing before the [MASK]. We take the average over all the sub-words for entities tokenized into sub-words. For a set of prompts , the CBS per entity type for an LM is computed by averaging over all . LMs are considered more Western-biased as its CBS gets closer to 100%.
@@ Appendix D Language Models Details
Lan et al. (2020): a bilingual English-Arabic BERT model that outperforms other multilingual models in zero-shot transfer from English to Arabic. GigaBERT666huggingface.co/lanwuwei/GigaBERT-v3-Arabic-and-English is trained on the Arabic and English Gigaword corpora, Arabic and English Wikipedia, and the OSCAR corpus. We also use a version of GigaBERT, referred to as GigaBERT-CS777huggingface.co/lanwuwei/GigaBERT-v4-Arabic-and-English, which is further pre-trained on Code-Switched data. Both models are in the base architecture.
@@
Scao et al. (2022): a 176 billion parameter multilingual LLM trained on 46 natural languages and 13 programming languages. The language-specific training data largely came from the OSCAR corpus Suárez et al. (2019). We used the HuggingFace inference API888https://huggingface.co/inference-api to prompt BLOOM which returns token log probabilities when using the details:true parameter.
@@
a 175 billion parameter LM. We experiment with OpenAI’s text-davinci-003 model which has been instruction fine-tuned. The data used to train the GPT-3.5 model has not been publicly disclosed. We retrieve token log probabilities using the OpenAI completions API endpoint999https://platform.openai.com/docs/api-reference/completions with the logprobs:1 parameter.
@@ Appendix E Pre-training Corpora Details
We use the September 2020 dump of Arabic Wikipedia101010https://dumps.wikimedia.org/ used in training AraBERT models Antoun et al. (2020).
@@ 1.5B Corpus:
News articles from the Lebanese Assafir newspaper111111https://en.wikipedia.org/wiki/As-Safir used in training AraGPT2 Antoun et al. (2021) and AraBERTv2 Antoun et al. (2020).
@@
The Open Super-large Crawled Almanach coRpus Suárez et al. (2019) is a multilingual partition of CommonCrawl121212https://commoncrawl.org/. We use the Arabic subset of the corpus.
@@ F.2.2 Experimental Details
We used a learning rate of 5e-5 and the AdamW Optimizer. We train models for 5 epochs and set the batch size to 8. Fine-tuning was performed on 1 NVIDIA A100 GPU. Since the HARD Elnagar et al. (2018) dataset for Arabic sentiment analysis is originally imbalanced in terms of sentiment labels, we took a random sample of 30k sentences from the dataset balanced across positive/negative/neutral sentiment for our experiment.

### 20 Mar 2024 v4

@@ Abstract
As the reach of large language models (LMs) expands globally, their ability to cater to diverse cultural contexts becomes crucial. Despite advancements in multilingual capabilities, models are not designed with appropriate cultural nuances. In this paper, we show that multilingual and Arabic monolingual LMs exhibit bias towards entities associated with Western culture. We introduce CAMeL, a novel resource of 628 naturally-occurring prompts and 20,368 entities spanning eight types that contrast Arab and Western cultures. CAMeL provides a foundation for measuring cultural biases in LMs through both extrinsic and intrinsic evaluations. Using CAMeL, we examine the cross-cultural performance in Arabic of 16 different LMs on tasks such as story generation, NER, and sentiment analysis, where we find concerning cases of stereotyping and cultural unfairness. We further test their text-infilling performance, revealing the incapability of appropriate adaptation to Arab cultural contexts. Finally, we analyze 6 Arabic pre-training corpora and find that commonly used sources such as Wikipedia may not be best suited to build culturally aware LMs, if used as they are without adjustment. We will make CAMeL publicly available at: https://github.com/tareknaous/camel
@@
We show that CAMeL entities and prompts enable cross-cultural testing of LMs in versatile experimental setups, including story generation, NER, sentiment analysis, and text infilling (§4). We benchmark 16 LMs pre-trained with Arabic data (§4.1). Our results reveal concerning cases of cultural stereotypes in LM-generated stories, such as the association of Arab names with poverty/traditionalism (§4.2), and cultural unfairness, such as better NER tagging performance of Western entities and higher association of Arab entities with negative sentiment (§4.3). We further show that LMs exhibit high levels of preference towards Western-associated entities even when prompted by contexts uniquely suited for Arab culture-associated entities (§4.4).
@@ 4.1 Language Models
We consider LMs that have been intentionally trained for Arabic. For monolingual LMs, we use AraBERT Antoun et al. (2020), ARBERT Abdul-Mageed et al. (2021), and CAMeLBERT Inoue et al. (2021); we compare CAMeLBERT to its variants trained exclusively on Dialectal Arabic (CAMeLBERT-DA) or Modern Standard Arabic (CAMeLBERT-MSA). Additionally, we use models trained on Arabic tweets such as MARBERT Abdul-Mageed et al. (2021) and AraBERT-T. We also include AraGPT2 Antoun et al. (2021). For multilingual LMs, besides mBERT, XLM-R Conneau et al. (2020), BLOOM Scao et al. (2022), GPT-3.5 and GPT-4, we use Arabic-English bilingual JAIS Sengupta et al. (2023), GigaBERT and GigaBERT-CS Lan et al. (2020), which was further trained on synthetic Arabic-English Code-Switched data. We also use AceGPT Huang et al. (2023), an instruction-tuned version of Llama2 Touvron et al. (2023) on localized Arabic instructions. Lastly, we use mT5 Xue et al. (2021) and its recent instruction-tuned counterpart Aya Üstün et al. (2024). We use the base () and large () versions whenever available. More details about all the LMs used can be found in Appendix D.
@@ 4.3 Fairness in NER and Sentiment Analysis
We create models capable of performing Arabic NER and sentiment prediction by fine-tuning LMs on datasets commonly used in Arabic NLU benchmarks Elmadany et al. (2023); Abdul-Mageed et al. (2021). We use the ANERCorp Benajiba et al. (2007) dataset for NER (name and location tags were used only) and HARD dataset Elnagar et al. (2018) for sentiment analysis. For GPT-type LMs, we perform in-context learning with 5-shot examples (see prompts in Appendix F.2).
@@ 4.4 Culturally-Appropriate Text Infilling
where is the LM’s probability of an entity filling the masked token. We evaluate LMs with BERT-type architecture using the full prompts with a [MASK] token for text-infilling and GPT-type/T5-type LMs using only the portion of the prompt appearing before the [MASK]. We take the average over all the sub-words for entities tokenized into sub-words. For a set of prompts , the CBS per entity type for an LM is computed by averaging over all . LMs are considered more Western-biased as its CBS gets closer to 100%.
@@ Appendix D Language Models Details
CAMeLBERT
Inoue et al. (2021): a BERT model trained on a variety of corpora that include Modern Standard Arabic, Dialectal Arabic, and Classical Arabic. We use the largest version of the model trained on 666https://huggingface.co/CAMeL-Lab/bert-base-arabic-camelbert-mix. We also compare to its variants: CAMeLBERT-DA777https://huggingface.co/CAMeL-Lab/bert-base-arabic-camelbert-da, which is trained only on Arabic dialects, and CAMeLBERT-MSA888https://huggingface.co/CAMeL-Lab/bert-base-arabic-camelbert-msa which is trained only on Modern Standard Arabic.
@@
Lan et al. (2020): a bilingual English-Arabic BERT model that outperforms other multilingual models in zero-shot transfer from English to Arabic. GigaBERT999huggingface.co/lanwuwei/GigaBERT-v3-Arabic-and-English is trained on the Arabic and English Gigaword corpora, Arabic and English Wikipedia, and the OSCAR corpus. We also use a version of GigaBERT, referred to as GigaBERT-CS101010huggingface.co/lanwuwei/GigaBERT-v4-Arabic-and-English, which is further pre-trained on Code-Switched data. Both models are in the base architecture.
@@
Scao et al. (2022): a 176 billion parameter multilingual LLM trained on 46 natural languages and 13 programming languages. The language-specific training data largely came from the OSCAR corpus Suárez et al. (2019). We used the HuggingFace inference API111111https://huggingface.co/inference-api to prompt BLOOM which returns token log probabilities when using the details:true parameter.
@@
AceGPT
mT5
AYA
Üstün et al. (2024): A 13 billion parameter model that performs instruction fine-tuning of mT5 in 101 languages to expand language coverage and improve performance in low-resource languages.
@@
a 175 billion parameter LM. We experiment with OpenAI’s text-davinci-003 model which has been instruction fine-tuned. The data used to train the GPT-3.5 model has not been publicly disclosed. We retrieve token log probabilities using the OpenAI completions API endpoint121212https://platform.openai.com/docs/api-reference/completions with the logprobs:1 parameter.
@@ Appendix E Pre-training Corpora Details
We use the September 2020 dump of Arabic Wikipedia131313https://dumps.wikimedia.org/ used in training AraBERT models Antoun et al. (2020).
@@ 1.5B Corpus:
News articles from the Lebanese Assafir newspaper141414https://en.wikipedia.org/wiki/As-Safir used in training AraGPT2 Antoun et al. (2021) and AraBERTv2 Antoun et al. (2020).
@@
The Open Super-large Crawled Almanach coRpus Suárez et al. (2019) is a multilingual partition of CommonCrawl151515https://commoncrawl.org/. We use the Arabic subset of the corpus.
@@ F.2.2 Experimental Details
We used a learning rate of 5e-5 and the AdamW Optimizer. We train models for 5 epochs and set the batch size to 8. Fine-tuning was performed on 1 NVIDIA A100 GPU. We fine-tuned Aya and mT5 using LoRA Hu et al. (2021) and 4-bit quantization. We set LoRa hyperparameters as follows: rank=8, alpha=16, dropout=0.05. Since the HARD Elnagar et al. (2018) dataset for Arabic sentiment analysis is originally imbalanced in terms of sentiment labels, we took a random sample of 30k sentences from the dataset balanced across positive/negative/neutral sentiment for our experiment.